Hello Math Wizards! I have the answer to the ff: problems but I want to confirm that my answer is correct.
Find the center, vertices and foci of the ellipse given by each equation.Sketch the graph.
1. x^2/49 + y^2/36 =1
2. ( x + 3 )^2/9 + ( y - 2)^2/16 = 1
Thanks,
Jo

Answer by stanbon(48569)

Find the center, vertices and foci of the ellipse given by each equation.Sketch the graph.
1. x^2/49 + y^2/36 =1
a=7, b=6
a^2=b^2+c^2
49=36+c^2
c=sqrt(13)
Center is at (0.0)
Vertices are at (7,0),(-7,0)
Foci are at (sqrt(13),0), (-sqrt(13),0)
2. ( x + 3 )^2/9 + ( y - 2)^2/16 = 1
Center is at (-3,2)
a=4,b=3
c^2=a^2-b^2
c^2=16-9=7
c=sqrt(7)
Vertices are at (-3,2-4), (-3,2+4)
Foci are at (-3,2-sqrt7) and (-3,2+sqrt7)
Cheers,
Stan H.